OpenStudy (calculusxy):

MEDAL!!! Target is having a sale in which every Sony TV is 35% off the regular price. Best Buy is having a sale in which $120 is deducted from the regular price of every Sony TV. Part A Josh is comparing the price of the same Sony TV in both stores. In both stores the TV normally sell for $415. Which store has the better bargain? Part B Susie buys three identical Sony TV at target. She pays $597.25 after the discount. What is the regular price of each TV? Round your answer to the nearest penny.

OpenStudy (calculusxy):

@mathmale

OpenStudy (mathmale):

What kind of help do you need? What is the regular price of this TV? If Target cuts that price by 35%, how much will a purchaser save on this TV?

OpenStudy (calculusxy):

For the first question, I put Target.

OpenStudy (mathmale):

Please discuss why you chose Target over Best Buy.

OpenStudy (calculusxy):

Okay, so for Target, I did: \[$415 \times 0.35 = $145.25\] then subtracted: \[$415 - $145 = $269.75\] For Best Buy: \[$415 - $120 = $295\] Then i decided that: \[$269. 75 < $295\]

OpenStudy (calculusxy):

@mathmale

OpenStudy (calculusxy):

@satellite73

OpenStudy (mathmale):

thanks for showing how you decided to choose to buy your TV from Target. i agree with you.

OpenStudy (calculusxy):

Can you help me with Part B? That's the one i need the most

OpenStudy (calculusxy):

I thought that it was $306.28

OpenStudy (mathmale):

Please share your thoughts about Part B, so that we can build upon what you already know.

OpenStudy (calculusxy):

okay

OpenStudy (mathmale):

Sounds like this lucky girl gets 3 TVs for only $597.25. Is that your interpretation, or do you think $597.25 is the sale price for EACH of the 3 TVs?

OpenStudy (calculusxy):

i did: 65% of ____ = 597.25 \[\frac{ is }{ of }= \frac{ % }{ 100 }\] i replaced those: \[\frac{ 597.25 }{ x } = \frac{ 13 }{ 20 }\] i then got: 11945=13x x = 918.846 ~ 918.85 since it is i believe that 918.85 is the total regular price, that's why divided by 3 and got 306.282 ~ $306.28

OpenStudy (calculusxy):

@mathmale

OpenStudy (calculusxy):

\[\frac{ is }{ of }=\frac{ percent }{ 100 }\]

OpenStudy (mathmale):

We could check your result as follows: Multiply your "regular price" figure by 0.65 and see whether or not that results in $597.25. If you do get that result, you can assume that your answer is correct.

OpenStudy (calculusxy):

Thanks @mathmale

OpenStudy (mathmale):

Your "regular price" amount is $918.85. Multiply that by 0.65: 0.65($918.85)=$597.25. What does this result tell you?